# Attractors for the Wave Equation with Nonlinear Energy Critical Sources and Dissipation

### v2

@@ 2.2 Statements of the Main Results
$\begin{aligned} &\int_{0}^{T}u_{tt}D_{\epsilon}udxdt+\int_{0}^{T}\int_{\Omega}\nabla u\nabla D_{\epsilon}udxdt\\ &+\int_{0}^{T}\int_{\Omega}g(u_{t})D_{\epsilon}udxdt+\int_{0}^{T}\int_{\Omega}f(u)D_{\epsilon}udxddt=\int_{0}^{T}\int_{\Omega}hD_{\epsilon}udxdt.\end{aligned}$ (3.38)
@@
$\lim_{\epsilon\rightarrow 0}\int_{0}^{T}u_{tt}D_{\epsilon}udxdt=\frac{1}{2}\|u_{t}(T)\|^{2}-\frac{1}{2}\|u_{t}(0)\|^{2}.$ (3.39)
@@
$\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\int_{\Omega}\nabla uD_{\epsilon}(\nabla u)dxdt=\frac{1}{2}\|\nabla u(T)\|^{2}-\frac{1}{2}\|\nabla u(0)\|^{2}.$ (3.40)
@@
$\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\int_{\Omega}g(u_{t})D_{\epsilon}udxdt=\int_{0}^{T}\int_{\Omega}g(u_{t})u_{t}dxdt.$ (3.41)
@@
$\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\int_{\Omega}f(u_{t})D_{\epsilon}udxddt=\int_{0}^{T}\int_{\Omega}f(u)u_{t}dxddt.$ (3.42)
@@
$\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\int_{\Omega}hD_{\epsilon}udxddt=\int_{0}^{T}\int_{\Omega}hu_{t}dxddt.$ (3.43)
@@
$\begin{aligned} &\frac{1}{2}\left[\,\|u_{t}(T)\|^{2}-\|u_{t}(0)\|^{2}\,\right]+\frac{1}{2}\left[\,\|\nabla u(T)\|^{2}-\|\nabla u(0)\|^{2}\,\right]\\ &+\int_{0}^{T}\int_{\Omega}g(u_{t})u_{t}dxddt+\int_{0}^{T}\int_{\Omega}f(u)u_{t}dxdt=\int_{0}^{T}\int_{\Omega}hu_{t}dxdt.\end{aligned}$ (3.44)
@@
$\int_{0}^{T}\int_{\Omega}f(u)u_{t}dxdt=\int_{\Omega}F(u(T))dx-\int_{\Omega}F(u(0))dx$
and
$\int_{0}^{T}\int_{\Omega}hu_{t}dxdt=\int_{\Omega}hu(T)dx-\int_{\Omega}hu(0)dx,$
@@
3.3 Proof of Theorem 2.2-(III): Hadamard Well-posedness
@@
$\begin{aligned} \int_{\Omega}\left[g(u^{1}_{t})-g(u^{2}_{t})\right]z_{t}dx=&\;\int_{\Omega}\int_{0}^{1}g^{\prime}(\theta u^{1}_{t}+(1-\theta)u^{2}_{t})d\theta|z_{t}|^{2}dx\\ \geq&\;\kappa_{0}\int_{\Omega}\int_{0}^{1}|\theta u^{1}_{t}+(1-\theta)u^{2}_{t}|^{4}d\theta|z_{t}|^{2}dx\geq 0.\end{aligned}$ (3.46)
@@
$\begin{aligned} \int_{\Omega}\left[f(u^{1})-f(u^{2})\,\right]z_{t}dx=&\;\frac{1}{2}\frac{d}{dt}\left[\int_{\Omega}\int_{0}^{1}f^{\prime}(\theta u^{1}+(1-\theta)u^{2})d\theta|z|^{2}dx\right]\\ &-\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\theta u^{1}+(1-\theta)u^{2})\left(\,\theta u^{1}_{t}+(1-\theta)u^{2}_{t}\,\right)d\theta|z|^{2}dx.\end{aligned}$ (3.47)
Then, using (3.46) and (3.47), it follows from (3.45) that
$\begin{aligned} \frac{d}{dt}E_{z}(t)\leq\underbrace{K_{f}\int_{\Omega}zz_{t}dx}_{\mathrm{I}_{1}}+\underbrace{\frac{1}{2}\int_{\Omega}\int_{0}^{1}\left|f^{\prime\prime}(\theta u^{1}+(1-\theta)u^{2})\right|\left|\,\theta u^{1}_{t}+(1-\theta)u^{2}_{t}\,\right|d\theta|z|^{2}dx}_{\mathrm{I}_{2}},\end{aligned}$ (3.48)
@@
  1. 1.
    The arguments given above provide uniqueness and continuous dependence in “one shot”. This should be contrasted with [5] where continuous dependence requires additional regularity of solutions.
  2. 2.
@@
3.4 Generation of a Dynamical System
Recall that a dynamical system is a pair of objects $(X,S_{t})$ consisting of a complet metric space $X$ and a family of continuous mappings $\{S_{t}:\,t\in\mathbb{R}^{+}\}$ of $X$ into itself with the semigroup properties:
$S_{0}=I,\quad S_{t+\tau}=S_{t}\circ S_{\tau}.$
$S_{t}$ is called evolution semigroup (or evolution operator). Moreover, $\{S_{t}:\,t\in\mathbb{R}^{+}\}$ is a strogly continuous semigroup of bounded linear operators on $X$ if
$\lim_{t\rightarrow 0}S_{t}x=x\quad\mbox{for every}\quad x\in X.$
It is also called: semigroup of class $C_{0}$ or simply a $C_{0}$-semigroup.
@@ 3.4 Generation of a Dynamical System
$\|S_{t}U_{0}^{1}-S_{t}U_{0}^{2}\|_{\mathcal{H}}^{2}\leq e^{c_{R}t}\|U_{0}^{1}-U_{0}^{2}\|_{\mathcal{H}}^{2},\quad\forall t\in[0,T],$ (3.52)
@@
4 Proof of Theorem 2.4: Global Attractors
The proof of the existence of a compact global tractor is established by the results obtained in Propositions 4.1 and 4.3 below. Proposition 4.1 shows that the dynamical system $(\mathcal{H},S_{t})$ is ultimately dissipative and Proposition 4.3 shows that $(\mathcal{H},S_{t})$ is asymptotically compact. Therefore, the existence of a compact global attractor follows as a direct application of Theorem 7.1 (in Appendix 7). Furthermore, the characterization of the attractor $\mathfrak{A}\equiv\mathcal{M}^{u}(\mathcal{N})$ is guaranteed by applying Theorem 7.2 using Proposition 4.4, demonstrated below, which guarantees that the dynamic system $(\mathcal{H},S_{t})$ is gradient. The details follow.
4.1 Ultimate Dissipativity
Proposition 4.1.
Proof.
Let $B$ be a bounded subset in $\mathcal{H}$ and let $U(t)=S_{t}U_{0}$ be a weak solution of problem (LABEL:P) with $U_{0}\in\mathcal{H}$. In order, let us consider the modified functional
$\mathcal{J}(U(t)):=E(U(t))+L,$
where $L:=\frac{1}{\omega\lambda_{1}}\|h\|^{2}+C_{\nu}|\Omega|$ and $E$ is the energy functional defined in (2.9). By argument similar to those used in (3.8), we obtain
[…]
Using the energy equality (2.8) with $s=t$ and $t=t+1$, we obtain
$\int_{t}^{t+1}\int_{\Omega}g(u_{t})u_{t}dxds=E(U(t))-E(U(t+1))=\mathcal{J}(U(t))-\mathcal{J}(U(t+1))=:\mathcal{D}(t).$
From Assumption (2.1), we get
$\int_{t}^{t+1}\|u_{t}(s)\|^{6}_{6}ds\leq\frac{1}{\kappa_{0}}\mathcal{D}(t).$ (4.2)
From Hölder inequality with $\frac{2}{3}+\frac{1}{3}=1$, it follows from (4.2) that
$\int_{t}^{t+1}\|u_{t}(s)\|^{2}ds\leq|\Omega|^{2/3}\left(\int_{t}^{t+1}\|u_{t}(s)\|^{6}_{6}ds\right)^{1/3}\leq\frac{|\Omega|^{2/3}}{\kappa_{0}^{1/3}}\mathcal{D}(t)^{1/3}.$ (4.3)
$\|u_{t}(t_{j})\|^{2}\leq 4\int_{t}^{t+1}\|u_{t}(s)\|^{2}ds\leq\frac{4|\Omega|^{2/3}}{\kappa_{0}^{1/3}}\mathcal{D}(t)^{1/3},\quad j=1,2.$ (4.4)
$\begin{aligned} \int_{t_{1}}^{t_{2}}E(s)ds+\frac{1}{2}\int_{t_{1}}^{t_{2}}\|\nabla u(s)\|^{2}ds=&\;\frac{3}{2}\int_{t_{1}}^{t_{2}}\|u_{t}(s)\|^{2}ds-\int_{t_{1}}^{t_{2}}\int_{\Omega}g(u_{t})udxds\\ &\;+\int_{\Omega}u(t_{1})u_{t}(t_{1})dx-\int_{\Omega}u(t_{2})u_{t}(t_{2})dx\\ &\;+\int_{t_{1}}^{t_{2}}\left[\int_{\Omega}F(u)dx-\int_{\Omega}f(u)udx\right]ds.\end{aligned}$ (4.5)
From Assumption (2.5), we get
$\int_{t_{1}}^{t_{2}}\left[\int_{\Omega}F(u)dx-\int_{\Omega}f(u)udx\,\right]ds\leq\frac{\nu}{2}\int_{t_{1}}^{t_{2}}\int_{\Omega}|u|^{2}dxds\leq\frac{\nu}{2\lambda_{1}}\int_{t_{1}}^{t_{2}}\|\nabla u(s)\|^{2}ds.$ (4.6)
$\begin{aligned} \int_{t_{1}}^{t_{2}}\mathcal{J}(U(s))ds+\frac{\omega}{2}\int_{t_{1}}^{t_{2}}\|\nabla u(s)\|^{2}ds\leq&\;\frac{3}{2}\int_{t_{1}}^{t_{2}}\|u_{t}(s)\|^{2}ds-\int_{t_{1}}^{t_{2}}\int_{\Omega}g(u_{t})udxds\\ &\;+\int_{\Omega}u(t_{1})u_{t}(t_{1})dx-\int_{\Omega}u(t_{2})u_{t}(t_{2})dx+L.\end{aligned}$ (4.7)
$-\int_{t_{1}}^{t_{2}}\int_{\Omega}g(u_{t})udxds$ $\leq$ $\kappa_{1}\int_{t_{1}}^{t_{2}}\int_{\Omega}(1+|u_{t}|^{4})|u_{t}||u|dxds$
$\leq$ $\kappa_{1}\left(\int_{t}^{t+1}\|u_{t}(s)\|^{2}ds\right)^{1/2}\left(\int_{t}^{t+1}\|u(s)\|^{2}ds\right)^{1/2}$
$+\kappa_{1}\left(\int_{t_{1}}^{t_{2}}\|u_{t}(s)\|^{6}_{6}ds\right)^{5/6}\left(\int_{t}^{t+1}\|u(s)\|^{6}_{6}ds\right)^{1/6}$
$\leq$ $\frac{\kappa_{1}}{\lambda_{1}^{1/2}}\left(\int_{t}^{t+1}\|u_{t}(s)\|^{2}ds\right)^{1/2}\left(\int_{t}^{t+1}\|\nabla u(s)\|^{2}ds\right)^{1/2}$
$+\frac{\kappa_{1}}{\lambda^{1/2}}\left(\int_{t_{1}}^{t_{2}}\|u_{t}(s)\|^{6}_{6}ds\right)^{5/6}\left(\int_{t}^{t+1}\|\nabla u(s)\|^{6}ds\right)^{1/6}$
$\leq$ $\frac{2\kappa_{1}}{(\omega\lambda_{1})^{1/2}}\left(\frac{|\Omega|^{2/3}}{\kappa_{0}^{1/3}}\mathcal{D}(t)^{1/3}\right)^{1/2}\sup_{t\leq s\leq t+1}\mathcal{J}(U(s))^{1/2}$
$+\frac{2\kappa_{1}}{(\omega\lambda)^{1/2}}\left(\frac{1}{\kappa_{0}}\mathcal{D}(t)\right)^{5/6}\sup_{t\leq s\leq t+1}\mathcal{J}(U(s))^{1/2}$
From embedding $H^{1}\hookrightarrow H^{0}$ and (4.4), we get
$\int_{\Omega}u(t_{1})u_{t}(t_{1})dx-\int_{\Omega}u(t_{2})u_{t}(t_{2})dx$ $\leq$ $\|u(t_{1})\|\|u_{t}(t_{1})\|+\|u(t_{2})\|\|u_{t}(t_{2})\|$
$\leq$ $\frac{1}{\lambda_{1}^{1/2}}\left[\,\|\nabla u(t_{1})\|\|u_{t}(t_{1})\|+\|\nabla u(t_{2})\|\|u_{t}(t_{2})\|\,\right]$
$\leq$ $\frac{4|\Omega|^{1/3}}{\kappa_{0}^{1/6}\lambda_{1}^{1/2}}\mathcal{D}(t)^{1/6}\sup_{t\leq s\leq t+1}\|\nabla u(s)\|$
$\leq$ $\frac{16|\Omega|^{2/3}}{\varepsilon\omega\kappa_{0}^{1/3}\lambda_{1}}\mathcal{D}(t)^{1/3}+\varepsilon\sup_{t\leq s\leq t+1}\mathcal{J}(U(t)).$
Returning to (4.5), we obtain
$\int_{t_{1}}^{t_{2}}\mathcal{J}(U(s))ds$ $\leq$ $C_{0}\mathcal{D}(t)^{1/3}+C_{1}\mathcal{D}(t)^{5/3}+2\varepsilon\sup_{t\leq s\leq t+1}\mathcal{J}(U(s))+L,$ (4.8)
$\mathcal{J}(U(t+1))\leq\mathcal{J}(U(\tau))=\frac{1}{t_{2}-t_{1}}\int_{t_{1}}^{t_{2}}\mathcal{J}(U(s))ds\leq 2\int_{t_{1}}^{t_{2}}\mathcal{J}(U(s))ds.$ (4.9)
Hence, from (4.2), (4.9), and (4.8), we get
$\mathcal{J}(U(t))\leq\left[\,4C_{0}+4C_{1}\mathcal{D}(t)^{4/3}+2\mathcal{D}(t)^{2/3}\,\right]\mathcal{D}(t)^{1/3}+4L.$
From the last inequality and using that $(a+b)^{3}\leq 4(a^{3}+b^{3})$, we have
$\mathcal{J}(U(t))^{3}\leq 4\left[\,4C_{0}+4C_{1}\mathcal{D}(t)^{4/3}+2\mathcal{D}(t)^{2/3}\,\right]^{3}\left[\,\mathcal{J}(U(t))-\mathcal{J}(U(t+1))\,\right]+[2^{8/3}L]^{3}.$ (4.10)
Using the definition of $\mathcal{D}(t)$ and the energy equality (2.8) we obtain that $|\mathcal{D}(t)|\leq 2|E(U_{0})|$. Then, we infer that
Thus, (4.10) can be rewritten as
$\mathcal{J}(U(t))^{3}\leq C_{0}\left[\,\mathcal{J}(U(t))-\mathcal{J}(U(t+1))\,\right]+\left[2^{8/3}L\right]^{3}.$ (4.11)
Applying Nakao’s lemma [26, Lemma 2.1] to (4.11), we conclude that
$\mathcal{J}(U(t))\leq\left(2C_{0}^{-1}(t-1)^{+}+\mathcal{J}(U(0))^{-2}\right)^{-\frac{1}{2}}+2^{8/3}L.$ (4.12)
Using (4.1), it follows from (4.12) that
[…]
and
[…]
Then, choosing $R^{2}>\frac{2^{14/3}L}{\omega}$, we conclude, from (4.13) and (4.14) that the ball
[…]
is an absorbing set in $\mathcal{H}$, for the dynamical system $(\mathcal{H},S_{t})$. ∎
@@ 4.1 Ultimate Dissipativity
In the particular case, when $h\equiv 0$ and $C_{\nu}=0$ in (2.5) we have that $\mathcal{J}(U(t))=E(U(t))$ and $L=0$. Thus, in (4.14) we obtain that
[…]
That is, the dynamical system $(\mathcal{H},S_{t})$ possesses, in its natural topological phase space $\mathcal{H}$ a compact global attractor $\mathfrak{A}=\{(0,0)\in\mathcal{H}\}$ (the set of a single point).
4.2 Asymptotic Compactness
@@
From Theorem 2.2-$\bf(IV)$} every weak solution $U(t)=(u(t),u_{t}(t))$ of Eq. (LABEL:P) satisfies the energy equality:
$E(U(t))+\int_{0}^{t}\int_{\Omega}g(u_{t})u_{t}dxds=E(U(0)),$ (4.15)
where
[…]
Consider the standard perturbed Lyapunov’s function $V$ with a perturbation $\epsilon$ given by
@@
$\varphi_{j}(t_{j})\rightharpoonup\lambda,\quad\lambda\in\mathcal{H}$ (4.17)
@@
$\lim\sup_{j\rightarrow\infty}V(\varphi_{j}(t_{j}))\leq V(\lambda)\leq\lim\inf_{j\rightarrow\infty}V(\varphi_{j}(t_{j})).$ (4.18)
@@
$-\int_{0}^{T}e^{-\epsilon(T-t)}M(U)dt$ $\leq$ $\lim\inf_{j}\int_{0}^{T}e^{-\epsilon(T-t)}(-M(U^{j}))dt$
$=$ $-\lim\sup_{j}\int_{0}^{T}e^{-\epsilon(T-t)}M(U^{j})dt.$

### v4

@@ 2.2. Statements of the Main Results
(3.30) $\begin{aligned} &\int_{0}^{T}u_{tt}D_{\epsilon}udxdt+\int_{0}^{T}\int_{\Omega}\nabla u\nabla D_{\epsilon}udxdt\\ &+\int_{0}^{T}\int_{\Omega}g(u_{t})D_{\epsilon}udxdt+\int_{0}^{T}\int_{\Omega}f(u)D_{\epsilon}udxddt=\int_{0}^{T}\int_{\Omega}hD_{\epsilon}udxdt.\end{aligned}$
@@
(3.31) $\lim_{\epsilon\rightarrow 0}\int_{0}^{T}u_{tt}D_{\epsilon}udxdt=\frac{1}{2}\|u_{t}(T)\|^{2}-\frac{1}{2}\|u_{t}(0)\|^{2}.$
@@
(3.32) $\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\int_{\Omega}\nabla uD_{\epsilon}(\nabla u)dxdt=\frac{1}{2}\|\nabla u(T)\|^{2}-\frac{1}{2}\|\nabla u(0)\|^{2}.$
@@
(3.33) $\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\int_{\Omega}g(u_{t})D_{\epsilon}udxdt=\int_{0}^{T}\int_{\Omega}g(u_{t})u_{t}dxdt.$
@@
(3.34) $\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\int_{\Omega}f(u_{t})D_{\epsilon}udxddt=\int_{0}^{T}\int_{\Omega}f(u)u_{t}dxddt.$
@@
(3.35) $\lim_{\epsilon\rightarrow 0}\int_{0}^{T}\int_{\Omega}hD_{\epsilon}udxddt=\int_{0}^{T}\int_{\Omega}hu_{t}dxddt.$
@@
(3.36) $\begin{aligned} &\frac{1}{2}\left[\,\|u_{t}(T)\|^{2}-\|u_{t}(0)\|^{2}\,\right]+\frac{1}{2}\left[\,\|\nabla u(T)\|^{2}-\|\nabla u(0)\|^{2}\,\right]\\ &+\int_{0}^{T}\int_{\Omega}g(u_{t})u_{t}dxddt+\int_{0}^{T}\int_{\Omega}f(u)u_{t}dxdt=\int_{0}^{T}\int_{\Omega}hu_{t}dxdt.\end{aligned}$
@@
$\int_{0}^{T}\int_{\Omega}f(u)u_{t}dxdt=\int_{\Omega}F(u(T))dx-\int_{\Omega}F(u(0))dx;~~~\int_{0}^{T}\int_{\Omega}hu_{t}dxdt=\int_{\Omega}hu(T)dx-\int_{\Omega}hu(0)dx,$
@@
3.3. Proof of Theorem 2.2-(III): Hadamard Well-posedness
@@
(3.38) $\begin{aligned} \int_{\Omega}\left[g(u^{1}_{t})-g(u^{2}_{t})\right]z_{t}dx=&\;\int_{\Omega}\int_{0}^{1}g^{\prime}(\theta u^{1}_{t}+(1-\theta)u^{2}_{t})d\theta|z_{t}|^{2}dx\\ \geq&\;\kappa_{0}\int_{\Omega}\int_{0}^{1}|\theta u^{1}_{t}+(1-\theta)u^{2}_{t}|^{4}d\theta|z_{t}|^{2}dx\geq 0.\end{aligned}$
@@
(3.39) $\int_{\Omega}\left[f(u^{1})-f(u^{2})\,\right]z_{t}dx=\;\frac{1}{2}\frac{d}{dt}\left[\int_{\Omega}\int_{0}^{1}f^{\prime}(\xi_{\theta})d\theta|z|^{2}dx\right]-\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\xi_{\theta})(\xi_{\theta})_{t}d\theta|z|^{2}dx.$
where $\xi_{\theta}=\theta u^{1}+(1-\theta)u^{2}$. Then, using (3.38) and (3.39), it follows from (3.37) that
(3.40) $\begin{aligned} \frac{d}{dt}E_{z}(t)\leq\underbrace{K_{f}\int_{\Omega}zz_{t}dx}_{\mathrm{I}_{1}}+\underbrace{\frac{1}{2}\int_{\Omega}\int_{0}^{1}\left|f^{\prime\prime}(\xi_{\theta})\right|\left|\,(\xi_{\theta})_{t}\,\right|d\theta|z|^{2}dx}_{\mathrm{I}_{2}},\end{aligned}$
@@
  1. (1)
    The arguments given above provide uniqueness and continuous dependence in “one shot”. This should be contrasted with [5] where continuous dependence requires an additional regularity of solutions.
  2. (2)
@@
3.4. Generation of a Dynamical System
@@
(3.44) $\|S_{t}U_{0}^{1}-S_{t}U_{0}^{2}\|_{\mathcal{H}}^{2}\leq e^{c_{R}t}\|U_{0}^{1}-U_{0}^{2}\|_{\mathcal{H}}^{2},\quad\forall t\in[0,T],$
@@
4. Proof of Theorem 2.4: Global Attractors
The existence of a compact global attractor is established by the results presented in Propositions 4.2, 4.3, 4.4, and 4.5. Proposition 4.2 shows that the dynamical system $(\mathcal{H},S_{t})$ is asymptotically compact, while Proposition 4.3 ensures that $(\mathcal{H},S_{t})$ is gradient. From Proposition 4.4, $\Phi(U)=E(U)$ is a Lyapunov function bounded from above on any bounded set of $\mathcal{H}$ and the set $\Phi_{R}=\{U:\Phi(U)\leq R\}$ is bounded for every $R$. Finally, it follows from Proposition 4.5 that the set $\mathcal{N}$ is bounded and there exists an absorbing ball for the system Corollary 2.29 [15]. Therefore, it follows directly from [16, Corollary 7.5.7] that the system $(\mathcal{H},S_{t})$ has a compact global attractor
$\mathfrak{A}=\mathrm{M}^{u}(\mathcal{N}).$
In particular, for the case $C_{\nu}=0$ and $h\equiv 0$ the attractor is the trivial $\mathfrak{A}=\{(0,0)\}.$
@@
In Appendix 7, we present a direct construction of an absorbing ball, which also provides specific parameters delineating the construct.
4.1. Asymptotic Compactness
@@
From Theorem 2.2-$\bf(IV)$} every weak solution $U(t)=(u(t),u_{t}(t))$ of Eq. (LABEL:P) satisfies the energy equality (2.8). Consider the standard perturbed Lyapunov’s function $V$ with a perturbation $\epsilon$ given by
@@
(4.2) $\varphi_{j}(t_{j})\rightharpoonup\lambda,\quad\lambda\in\mathcal{H}$
@@
(4.3) $\lim\sup_{j\rightarrow\infty}V(\varphi_{j}(t_{j}))\leq V(\lambda)\leq\lim\inf_{j\rightarrow\infty}V(\varphi_{j}(t_{j})).$
@@
$-\int_{0}^{T}e^{-\epsilon(T-t)}M(U)dt$ $\leq$ $\lim\inf_{j}\int_{0}^{T}e^{-\epsilon(T-t)}(-M(U^{j}))dt=-\lim\sup_{j}\int_{0}^{T}e^{-\epsilon(T-t)}M(U^{j})dt.$